LBC Preparations for Partner Science:
UT 2010 Jun 05

Observers: OKuhn, DThompson, MPedani (LBTO)
Telescope Operator: SAllanson
Instrument Support: JMorris

Summary

  • Testing M3 position with Lucifer; more tests needed tomorrow!
  • IA=-393, no change with respect to Default model
  • LBCR image with clear banding: lbcr.20100605.064931
  • difficulty restoring tick marks on LBC User Interface
    • uncheck one channel or the mirror, but have trouble getting the check marks to 'stick' afterwards ---

Details

Weather at begin of Night: T=13.5, RH=21%, Wind=4m/s

02:30 - Testing M3 position with Lucifer; M3 is supposed to be tilted

05:15 - End Testing with Lucifer; had some problems with guiding also; M3 was put to nominal position

05:16 We switch to LBC

05:45 Slew to WT10_313 to check Pointing at El=+85. DOFPIA also. We have some crashes; pupils are pretty bad, Mirrors ar 2^C colder than outsde air

Look at images:
lbcb.20100605.055910 & 055650 and lbcr.20100605.055907 & 060545

06:10 Still struggling with pupils; DOFPIA crashes at Blue. Manual intervention needed

06:45 DOFPIA converged and pointing adjusted. IA=-393.0, CA=-72.77 IE=-25.95

06:47 Slew to WT10_281 close to Sci Field to Correct Pointing & Co-Pointing

06:50 T=15.3, RH=20%, Wind=2.0m/s, T_Mirror=14.0

06:50 lbcr.20100605.064931 shows banding! (Issue track 1191 Note Added)

07:10 Slew to Sci Field SN2009DC; problems with Collimation. The OB seems not to read Filter Focus offsets but it's still getting -0.8mm from collimation procedure. We get Donuts!

07:35 It seems that TCS is not applying filter focus corrections and we have to apply them manually as global offsets but images are pretty good

08:19 DOFPIA on target. The problem was unticking Mirror on the OB GUI frown, sad smile

08:30 Restarted LBC and DOFPIA again on target

08:37 Again Run Sci OB on SN2009DC; Red Images ALL with FWHM ~2.5 pixels. So the Tech Chip IA seems to work fine

09:08 Slew to M16 for Pretty Picture and DOFPIA

09:17 Run M16 OB with gri filters PA=0.0; FWHM ~2.5 pixels; Images look very good

09:41 DOFPIA on same Field

09:43 T=15.0 RH=25%, Wind=2m/s T_mirror=14.6

09:59 Run M16 OB with gri filters PA=90.0

10:24 We take a short Pointing Model with LBC; started pointing log: 20100605.ptlog

ACT_0407 Pos#2
WT10_401 Pos#3
ACT_0345 Pos#4
WT10_281 Pos #5
WT10_196 Pos #6
WT10_234 Pos#7
WT10_096 Pos #8
WT10_168 Pos#9
ACT_0225 Pos#11
ACT_0325 Pos#12
ACT_0338 Pos#13
WT10_293 Pos#14
WT10_259 Pos#15
WT10_289 Pos#16
WT10_373 Pos#1
WT10_274 Pos#10

11:26 Slew to Flat Field

  • five g (B) and i (R) flats
  • five r (B) and r(R) flats
  • five B (B) flats ... missed the I-BESSEL flats as these were saturated. Five I-BESSEL sky flats were obtained in evening twilight the next night (20100606 UT).

In the next science run, I-BESSEL flats may be obtained. Or - we'll try to leave LBC-Red in the beam and take these sky flats tomorrow evening. For now, there are I-BESSEL SkyFlats available from as recently as 20100507 and 20100416.

11:59:00 10 biases with telescope stationary and lights off in dome.

-- Main. Marco Pedani - 05 Jun 2010
Topic revision: r4 - 06 Jun 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback